# Some models are useful, but for how long?: A decision theoretic approach to choosing when to refit large-scale prediction models

### v1

@@ Abstract
Large-scale prediction models (typically using tools from artificial intelligence, AI, or machine learning, ML) are increasingly ubiquitous across a variety of industries and scientific domains. Such methods are often paired with detailed data from sources such as electronic health records, wearable sensors, and omics data (high-throughput technology used to understand biology). Despite their utility, implementing AI and ML tools at the scale necessary to work with this data introduces two major challenges. First, it can cost tens of thousands of dollars to train a modern AI/ML model at scale. Second, once the model is trained, its predictions may become less relevant as patient and provider behavior change, and predictions made for one geographical area may be less accurate for another. These two challenges raise a fundamental question: how often should you refit the AI/ML model to optimally trade-off between cost and relevance? Our work provides a framework for making decisions about when to refit AI/ML models when the goal is to maintain valid statistical inference (e.g. estimating a treatment effect in a clinical trial). Drawing on portfolio optimization theory, we treat the decision of recalibrating versus refitting the model as a choice between “investing” in one of two “assets.” One asset, recalibrating the model based on another model, is quick and relatively inexpensive but bears uncertainty from sampling and the possibility that the other model is not relevant to current circumstances. The other asset, refitting the model, is costly but removes the irrelevance concern (though not the risk of sampling error). We explore the balancing act between these two potential investments in this paper.
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI) and machine learning (ML), large-scale models are becoming standard tools in scientific research and industry settings to to generate predictions of hard-to-measure outcomes [47]. For example, AI/ML techniques applied to electronic health records can identify metastatic cancers [56, 3], determine causes of death [15], or provide targeted real-time early warning systems for adverse health events such as sepsis in emergent settings [53]. Transformer models such as AlphaFold can predict protein structures from easily obtained amino acid sequences [22]. Deep learning of spatio-temporal dependencies in energy consumption data can forecast future consumption patterns [41]. Today, researchers and practitioners benefit from large generative AI models, which have afforded greater access to AI/ML infrastructure and have enabled the performance of more general tasks with a high degree of accuracy [55, 42]. Particularly in research, AI/ML predicted outcomes are then used in place of the true, unmeasured outcomes for downstream inference [i.e., finding associations, estimating a treatment effect; see 48, 6, 32].
The economic cost of computing to train these complex prediction models can easily reach into the tens of thousands of dollars [40]. There are, of course, organizations that exploit economies of scale and train general models for a wide variety of prediction tasks [30, 31], although even in these settings, some local training is typically required for high-quality predictions. Models also decay with time – user behavior changes, healthcare technology and norms adapt, physical systems evolve , and concept and model drift occur [33, 45, 49, 26, 57] – meaning that users must answer a fundamental question: what is the optimal strategy for maintaining model accuracy and relevance?
We address this question in the context of statistical inference, where a user wants to perform downstream statistical tests based on outcomes generated from said prediction model. Inference is critical in a range of scientific settings to learn associations between observable covariates and outcomes or estimate treatment effects in randomized experiments. For example, global public health researchers can estimate cause-specific death rates by regional differences in demographics using NLP-derived cause-of-death labels from verbal autopsies [15]. In the context of pragmatic clinical trials, trialists predict outcomes that may have occurred between regularly scheduled clinic visits to better estimate the real-world effectiveness of interventions in everyday clinical practice [16, 52].
@@
In this work, we propose a pragmatic strategy for choosing when to refit complex prediction models, when the goal is inference, by exploiting a novel connection to portfolio optimization theory. We compare three distinct strategies: (1) fully refitting the ML/AI model from scratch, (2) partially recalibrating predictions using a small new sample, or (3) simply retaining the existing model without update. The decision of whether to refit, recalibrate, or retain involves significant financial considerations. Fully refitting a large-scale predictive model can incur substantial computational resource, labor, and time costs [40]. Recalibrating uses a few labeled samples to adjust for bias induced by the prediction model, though we do not know until after making the decision how effective this will be. Instead, we only know how well a similar procedure worked in a previous calibration period. Retaining the existing model without updates, while cost-free in the short term, risks a gradual decline in accuracy and relevance of the model over time [46].
@@
We approach refitting versus recalibrating using portfolio optimization theory, where this decision is treated as ‘investing’ in one of two options, or ‘assets.’ This powerful connection, to our knowledge, has not yet been exploited in the machine learning literature. Portfolio theory comes with a natural language for expressing uncertainty. Refitting the model has statistical uncertainty, or risk. We know approximately how well our prediction algorithm will do based on prior fits of the model, but we expect some variation based on the inherent stochasticity of the prediction task. Recalibrating also introduces ambiguity, meaning that we do not know how much the prediction model has decayed from the last time it was trained and, thus, how effective our adjustment will be. Portfolio optimization theory has a similar conundrum, where an investor must choose between assets (stocks, bonds, etc.) that are volatile, but in a mostly predictable way, and more uncertain assets where the distribution of volatility is harder to predict. In both cases, it is impossible to characterize the optimal decision solely in terms of statistical properties – the user has incomplete information in a stochastic environment, which necessitates introducing idiosyncratic user preferences to avoid either risk or ambiguity.
We formalize this connection to portfolio optimization, after providing a full description of the three options, in Section 2. Decisions depend on a specified utility function that weighs the loss associated with each choice, plus user-specific preferences that represent aversion to risk and ambiguity. Specifying these preference parameters typically requires an extensive solicitation process. In Section 2.2, we propose an alternative that requires specifying only a Type 1 error rate. Section 3 provides simulation results and Section 4 gives an example decision process in the context of electric consumption forecasting. We conclude with limitations and future directions.
2 Determining the Optimal Strategy
2.1 Decision Rule
We assume that the user starts with a trained AI/ML model and access to a stable (calibration) period where they can collect new training feature (predictor) and ground-truth label (outcome) data. The goal is optimally decide how to maintain the model at some point in the future when its performance may have changed by an unknown amount, having seen the predictors and predicted outcomes leading up to that point. The three potential strategies trade-off between cost, risk, and ambiguity:
@@ 2.1 Decision Rule
    Retain the model - Cheap but uncertain. Responses may no longer be accurate/useful.
@@
    Refit the model - Expensive but certain. Need to collect new training data, plus pay model fitting cost. Inherent/natural variation in data generating process (risk).
@@
    Recalibrate the model - Intermediate option. Use a few labeled cases to estimate adjustment factor that accounts for concept drift. Natural variation in data generating process (risk) and uncertain quality of the correction factor (ambiguity).
As we only collect new data after the decision to refit or recalibrate, our strategy must reflect our present uncertainty surrounding the future model’s performance. To accomplish this, we exploit a novel parallel to selecting the optimal asset distribution in the presence of stochastic returns from portfolio optimization theory  [8, 29]. In choosing an asset distribution, the goal is to assign a fixed amount of wealth between assets, with stochastic returns, to maximize a given economic utility function. Our ‘assets’ are the three strategies, and we seek to minimize the mean squared error (MSE) in statistical inference subject cost, budget, and preferences over risk/ambiguity. Importantly, the MSE is for estimating the parameters of a model we wish to draw inference on, using outcomes predicted from our AI/ML model, and not the mean squared error of the AI/ML model itself. To formalize this, assume the maintainer of the AI/ML model is a rational actor [see 34]. Following Maccheroni et al. [27], we define the utility function, $U(\cdot)$, such that
$U(MSE,\lambda,\theta)=-1\times E_{P}(MSE)-\frac{\lambda}{2}\sigma^{2}_{P}(MSE)-\frac{\theta}{2}\sigma^{2}_{\mu}(MSE),$ (1)
where $E_{P}(MSE)$ is the expected MSE of the downstream inferential model, $\sigma_{P}^{2}(MSE)$ and $\sigma_{\mu}^{2}(MSE)$ are the variation in the MSE that are attributable to sampling (i.e., risk) versus variability in estimating the recalibration factor (i.e., ambiguity), respectively. The parameters $\lambda$ and $\theta$ are positive real numbers that indicate how averse one is to risk and ambiguity, respectively. Letting $MSE_{rec}$, $MSE_{ref}$, and $MSE_{ret}$ be the respective MSEs of the downstream inferential model after recalibrating, refitting, or retaining the AI/ML model, the optimal allocation of our budget between $MSE_{rec}$ and $MSE_{ref}$ is determined by allocations $w_{rec}$ and $w_{ref}$ such that
$w_{rec}\times MSE_{rec}+w_{ref}\times MSE_{rec}$
maximizes our utility. Differentiating Eq. (1) with respect to $\lambda$ and $\theta$ yields the system of equations
@@
where $\sigma_{P}(MSE_{ref},MSE_{rec})$ is the risk covariance for recalibrate versus refit. The solution is
@@
Finally, we choose the option has the larger allocation if it is better than retaining, which yields the final decision rule
Decision Rule={ $\text{Decision Rule}=\begin{cases}\text{ Refit is Optimal}&\text{If }\hat{w}_{ref}>\hat{w}_{rec}\text{ and }A>0\\ \text{ Recalibrate is Optimal}&\text{If }\hat{w}_{ref}>\hat{w}_{rec}\text{ and}B>0\\ \text{ Retain is Optimal }&\text{If }A<0\text{ and }B<0\end{cases}.$ (2)
2.2 Risk and Ambiguity Aversion
The above decision rule (2) relies on user-specified preferences for risk ($\lambda$) and ambiguity ($\theta$). Soliciting such preferences is an active area of research, and there exists a large body of literature on the topic [18, 12, 11, 28]. Two promising approaches, as described in Bonhomme and Weidner [9], are a context-specific approach versus a hypothesis testing approach. The context-specific approach elicits informative choices of $\lambda$ and $\theta$ using prior knowledge of the type of misspecification. The hypothesis testing approach is most salient when a user lacks such domain knowledge.
We develop a strategy based on the hypothesis testing approach. Immediately after AI/ML model training, we expect minimal drift, and thus retaining is likely the best option. As we are dealing with a stochastic decision-making process, there is always a nonzero probability that we could choose to refit or recalibrate during this period. We can decrease the probability that refitting or recalibrating is chosen by increasing $\lambda$ and $\theta$, our aversion parameters. We choose values of $\lambda$ and $\theta$ such that, early on, we have a less than $\alpha$ probability of choosing anything but to retain the current model by setting

### v2

# Some models are useful, but when?: A decision-theoretic approach to choosing when to refit large-scale prediction models

@@ Abstract
Large-scale prediction models using tools from artificial intelligence (AI) or machine learning (ML) are increasingly common across a variety of industries and scientific domains. Despite their effectiveness, training AI and ML tools at scale can cost tens or hundreds of thousands of dollars (or more); and even after a model is trained, substantial resources must be invested to keep models up-to-date. This paper presents a decision-theoretic framework for deciding when to refit an AI/ML model when the goal is to perform unbiased statistical inference using partially AI/ML-generated data. Drawing on portfolio optimization theory, we treat the decision of recalibrating a model or statistical inference versus refitting the model as a choice between “investing” in one of two “assets.” One asset, recalibrating the model based on another model, is quick and relatively inexpensive but bears uncertainty from sampling and may not be robust to model drift. The other asset, refitting the model, is costly but removes the drift concern (though not statistical uncertainty from sampling). We present a framework for balancing these two potential investments while preserving statistical validity. We evaluate the framework using simulation and data on electricity usage and predicting flu trends.
Keywords: Decision theory; Inference on Predicted Data; Arrow-Pratt Utility;
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI) and machine learning (ML), large-scale models are becoming standard tools in scientific research and industry settings to generate predictions of hard-to-measure outcomes. For example, AI/ML techniques applied to electronic health records can identify metastatic cancers (Yang et al., 2022), determine causes of death (Fan et al., 2024; Byass et al., 2019; McCormick et al., 2016), or provide targeted real-time early warning systems for adverse health events such as sepsis in emergent settings (Wong et al., 2021). Transformer models such as AlphaFold can predict protein structures from easily obtained amino acid sequences (Jumper et al., 2021). Deep learning of spatio-temporal dependencies in energy consumption data can forecast future consumption patterns  (Somu et al., 2021).
Typically, these predicted outcomes are not the final goal, but intermediate sources of auxiliary information to use when estimating a particular target of inference (e.g., a population statistic, regression coefficient, or treatment effect). For example, global public health researchers estimate how cause-of-death varies by demographics using cause-of-death predicted using NLP techniques narratives from caretakers (Fan et al., 2024). In pragmatic clinical trials, researchers predict outcomes that may have occurred between regularly scheduled clinic visits to estimate the effectiveness of interventions in everyday clinical practice  (Gamerman et al., 2019; Williams et al., 2015).
This estimation step faces statistical, sampling-based uncertainty. Since the outcomes are predicted, valid inference also depends on how well the prediction model performs. If the population for inference is the same as (or has a known relationship to) the population used to train the prediction model, then we can estimate, and correct for, the additional error introduced by the prediction model. This emerging area (Wang et al., 2020; Miao et al., 2023; Angelopoulos et al., 2023; Miao and Lu, 2024; Egami et al., 2024) is referred to in Hoffman et al. (2024) as “Inference on Predicted Data.” In many applied settings, however, the relationship between the population of interest for inference and the one used for training is opaque, and the practitioner must decide whether or not to use the model’s predictions without full information about this relationship. This paper addresses this increasingly ubiquitous conundrum. We present a decision-theoretic framework to decide when to refit an upstream prediction model when the goal is to perform unbiased downstream statistical inference on AI/ML-predicted outcomes. Drawing on portfolio optimization theory, we treat this problem as a choice between “investing” in one of two “assets.”
Performing inference using predicted outcomes requires managing traditional, sampling-based, statistical uncertainty as well as uncertainty about how appropriate the prediction model is to the current data setting. Mismatch between the mapping of features to outcomes between one population and another is known as “concept drift.” Concept drift can be gradual or sudden, making it difficult to quantify without extensive data from the new population. For example, patient behavior, consumer or voter preferences, and epidemiological context are highly volatile. In our empirical examples, we study a case where a policy change creates a rapid fluctuation in the cost of electricity and a case where changing epidemiological context and internet usage hamper the ability to predict flu trends. In both cases, concept drift renders inference using these predicted outcomes invalid.
The practitioner cannot eliminate uncertainty arising from concept drift without retraining the prediction model, which can cost upwards of tens of thousands of dollars or more and could be impractical in many settings. (Li et al., 2023; Sharir et al., 2020). The practitioner must therefore choose, using factors beyond the data, whether or not to trust the model. The decision relies on the practitioner’s preferences which depends on the consequences of making the wrong decision, which is inherently bespoke to a given situation. This could include the practitioner’s tolerance for making a mistake (and the situation-specific consequences thereof), the cost of model fitting and data collection, and the ratio of signal to noise in the parameter of interest.
@@
In this work, we propose a pragmatic strategy for choosing when to refit complex prediction models, when the goal is inference, by exploiting a novel connection to portfolio optimization theory. We compare three distinct strategies: (1) fully refitting the ML/AI model from scratch, (2) partially recalibrating predictions using a small new sample, or (3) simply retaining the existing model without update. The decision of whether to refit, recalibrate, or retain involves significant financial considerations. Fully refitting a large-scale predictive model can incur substantial computational resource, labor, and time costs (Smith, 2023). Recalibrating uses a few labeled samples to adjust for the bias induced by the prediction model, although we do not know until after making the decision how effective this will be. We only know how well a similar procedure worked in a previous calibration period. In situations where the concept drift is rapid, the recalibration will be inferior to a full refit. Retaining the existing model without updates, while cost-free in the short term, risks a gradual decline in accuracy and relevance of the model over time (Vela et al., 2022).
@@
We approach refitting versus recalibrating using portfolio optimization theory, where this decision is treated as ‘investing’ in one of two options, or ‘assets.’ This powerful connection, to our knowledge, has not yet been exploited in the machine learning or statistics literature. Portfolio theory comes with a natural language for expressing uncertainty and quantifying preferences. Refitting the model has statistical uncertainty, or risk. We know approximately how well our prediction algorithm will do based on prior fits of the model, but we expect some variation based on the inherent stochasticity of the prediction task. Recalibrating also introduces ambiguity, meaning that we do not know how much the prediction model has decayed from the last time it was trained and, thus, how effective our adjustment will be. Portfolio optimization theory has a similar conundrum, where an investor must choose between assets (stocks, bonds, etc.) that are volatile, but in a mostly predictable way, and more uncertain assets where the distribution of volatility is harder to predict. In both cases, it is impossible to characterize the optimal decision solely in terms of statistical properties – the user has incomplete information in a stochastic environment, which necessitates introducing idiosyncratic user preferences to avoid either risk or ambiguity.
The third option, retaining, we think of as the baseline choice. Retaining means that you believe the conditions that were used to fit your model now apply perfectly in your new setting (e.g. time or location). This is equivalent to assuming that your preferences are such that you have no aversion to ambiguity. As seen in our empirical application predicting flu trends, retaining your model is rarely reasonable in practice and can lead to misleading results. Retain is preferred if the cost of collecting even a small amount of data to recalibrate is extremely high, and the cost of retraining a model is also high. In such cases, the decision-making process is not interesting since the practitioner has no choices to make. Retaining, therefore, is a possibility under some circumstances but not a typical part of our decision-making paradigm. We will illustrate examples of the difficulties of choosing between these strategies in two data-driven examples. One will illustrate the issues in elasticity and electricity price estimation, while the latter will reassess the infamous Google Flu algorithm and evaluate when the algorithm should have been retrained.
Our work is part of a newly flourishing literature that emphasizes the role of the modeler (and their preferences and context) in making consequential decisions that will impact the interpretation of statistical results. The idea that a data scientist should transparently and reproducibly report all steps in the discovery pipeline is a key idea in Bin Yu’s “Veridical Data Science” (Yu and Kumbier, 2020). For example, quantifying the shared versus differing information learned from multiple models, and thus how likely one’s “best model” may have been chosen by random chance, has driven the study of the “Rashomon sets”(Rudin et al., 2024; Venkateswaran et al., 2024). Andrew Gelman brought to attention the connection between these decisions, which he calls “The Garden of Forking Paths,” to the subject of multiple comparisons and $p$-hacking (Gelman and Loken, 2019). Xiao-Li Meng described the effect of non-probabilistic decisions that one makes through his language on Design, Divine, and Device probability (Meng, 2022). In these and other works, much has been done to articulate, quantify, and describe how different choices at these decision points can lead to different models and potentially different conclusions. Our work, instead, focuses on systematically evaluating these options, weighing them against one’s preferences, and developing rules to make a decision in the face of this uncertainty.
2 Framework
2.1 Modeling Assumptions and Goals
We are motivated by studies with an expensive or hard-to-obtain outcome, $Y$, and potentially associated covariates $X$. The relationship between $Y$ and $X$ is governed by a stochastic, data generation process, $G$, which changes with respect to $t\in[0,T]$. : Y(t) = G(X, t). The practitioner’s goal is to describe the relationship between $Y$ and $X$ at time $t$ as parameterized by $\beta$ from interpretable model $\tilde{G}$. For instance, in political science, $Y$ could be a person’s political affiliation or voting patterns, which may require costly in-person interviews or exit surveys, $X$ could be various demographic variables such as age, sex, race, $G$ would be the true psycho-social-economic process by which these factors influences a person’s voting patterns, $\tilde{G}$, a logistic regression that a social scientist fits to model said the relationship, $\beta$, the coefficients of said logistic regression, and $t$ modeling how these preferences change over time. We use time to make our procedure concrete, but the method we propose is directly applicable in other continuous domains as well (e.g. distance).
Further, the practitioner has access to a pre-trained AI/ML model, $\hat{G}_{0}$, which predicts $Y$ using $X$, though they may not have access to the operating characteristics of this model or the data used to train it. In the previous example, this could be a natural language processing model that predicts voting behavior from online footprint Jiang et al. (2024); Sumi (2021). On one hand, such a model could potentially be a great boon. If the model provides reasonable estimates of $Y|X$ at a much cheaper cost, then one can eschew or cut down on the budget spent on collecting the hard-to-observe $Y$s. On the other hand, if the model is inaccurate or biased, then it is better to spend our budget mostly collecting real data and ignoring the predicted $Y$s. Further complicating the question, is the effect of concept drift. If the black-box model was fit in the past, then it is possible that it may have provided accurate estimates in the past, but is no longer useful for decisions in the future. In designing our decision-making framework, we would like our procedure to be able to take into account these various factors.
At a high level, this leaves the practitioner with three potential strategies with regard to how to use the said model in their discovery pipeline, while balancing the cost to collect $Y$ versus $X$, the risk versus and the ambiguity:
@@ 2.1 Modeling Assumptions and Goals
    Retain the model - collect $X$, then use the previously trained model to impute $Y$ to get $\tilde{G}_{0}(X)$ and regress $\tilde{G}_{0}(X)$ on $X$ to estimate $\beta$. - Cheap but uncertain. Responses may no longer be accurate/useful depending on model performance and the amount of concept drift.
@@
    Refit the model- Collect new $(X,Y)$ data to refit $\tilde{G}_{0}(X)$ the proceed as Retain.- Expensive but certain. Need to collect new training data, plus pay model fitting cost. Inherent/natural variation in data generating process (risk).
@@
    Recalibrate the model - Intermediate option. Use a few labeled cases to estimate the adjustment factor that accounts for concept drift. Natural variation in data generating process (risk) and uncertainty in estimating the correction factor (ambiguity).
In the succeeding sections, we will demonstrate how to build a decision-making framework for these three options and how to estimate the utility of each option.
2.2 Making the Decision
We approach the choice between retaining, refitting, or recalibrating as a decision theory problem where we are interested in making a decision in the presence of uncertainty. However, there are two sources of uncertainty: risk and ambiguity. Risk refers to the effect that the inherent stochasticity of the data generation process has on parameter estimation. Due to the stochastic nature of the data generation process, one would expect to observe a different value for Y if the experiment were rerun. With a different $Y$, regardless of whether one chooses a retain, refit, or recalibrate approach, one would get a slightly different estimate for $\beta$. This is the most common form of uncertainty that traditional frequentist hypothesis testing is built to quantify.
On the other hand, ambiguity refers to model uncertainty, or in other words, the uncertainty that arises from how uncertain one is about the true probabilistic model. For example, risk occurs when an actor is given perfect information about the data generation process $G$ but, due to natural stochasticity, chooses the wrong strategy. Ambiguity is when $G$ is not known and must be estimated, leading to a range of possible data generation models. Under this setup, decision-making is to be done without knowledge of the specification of the data-generating process or the AI/ML model, so both forms of uncertainty are present. In Section 2.3, we describe how to estimate both sources of uncertainty.
To build a decision framework that can handle risk and ambiguity, we employ a connection to a field that has well-studied the trade-offs of making decisions under different sources of uncertainty: portfolio optimization theory. Portfolio theory comes with a natural language for expressing uncertainty. One starts with a fixed budget for data collection. Then, one can treat choosing a strategy as ‘investing’ in an option, and after pursuing the said option, one can observe whether it provided the best return, in our case, gave us the best estimate of our parameter of interest in terms of mean squared error (MSE). Moreover, this gives us the ability to easily include fixed costs, such as the cost of training an AI/ML model, which is often a substantial burden in many studies. The MSE here is for estimating the parameters of a downstream model we wish to draw inference on using outcomes predicted from our upstream AI/ML model and not the mean squared error of the AI/ML model itself.
To choose an appropriate utility function for the decision-making process, one needs a utility function that places a high value on obtaining the best-performing result while penalizing the uncertainty in said result. In portfolio allocation, this is oft described as the trade-off between the expected return of an option versus its volatility. We choose as our objective function the Arrow-Pratt utility function, modified to penalize risk and ambiguity Maccheroni et al. (2013). We assume the person making the decision is a rational actor whose goal is to choose the strategy that maximizes an ambiguity-adjusted Arrow-Pratt utility function, $U(\cdot)$ of the form:
$U(MSE,\lambda,\theta)=-1\times E_{P}(MSE)-\frac{\lambda}{2}\sigma^{2}_{P}(MSE)-\frac{\theta}{2}\sigma^{2}_{\mu}(MSE).$ (1)
$E_{P}(MSE)$ is the expected MSE of our parameter of interest, $\beta$, $\sigma_{P}^{2}(MSE)$ and $\sigma_{\mu}^{2}(MSE)$ are the variation in the MSE that is attributable to sampling (ie, risk) versus variability in estimating the factor recalibration (ie, ambiguity), respectively. The parameters $\lambda$ and $\theta$ are positive real numbers that indicate how averse one is to risk and ambiguity, respectively, with higher values of $\lambda$ and $\theta$ corresponding to more aversion to risk and ambiguity, respectively. See Section 2.2.1 for a detailed discussion on how $\lambda$ and $\theta$ can be chosen in practice. Let $MSE_{rec}$, $MSE_{ref}$, and $MSE_{ret}$ be the MSEs of $\beta$ when taking each of three strategies of recalibrating, refitting, or retaining. Assuming that one does need to expend any additional funds when choosing to retain a model, the economic question becomes when given a fixed budget $c$, what is the optimal allocation of said budget to the recalibrate versus refitting strategies? Modeling $w_{rec}$ and $w_{ref}$ as weights such that $w_{rec}+w_{ref}=1$, we seek to solve for the weights such that the combined asset:
$w_{rec}\times MSE_{rec}+w_{ref}\times MSE_{ref}$
maximizes our utility function, Equation 1. Solving this yields a system of equations:
@@ 2.2 Making the Decision
where $\sigma_{P}(MSE_{ref},MSE_{rec})$ is the risk covariance for recalibrate versus refit. The solution occurs at:
@@
Finally, while a fractional allocation between strategies is a valid approach when buying financial assets, for choosing between data pipelines, we must make a discrete choice, so we utilize the following strategy to discretize our options:
Decision Rule={ $\text{Decision Rule}=\begin{cases}\text{ Refit is Optimal}&\text{If }\hat{w}_{ref}>\hat{w}_{rec}\text{ and }A>0\\ \text{ Recalibrate is Optimal}&\text{If }\hat{w}_{ref}<\hat{w}_{rec}\text{ and}B>0\\ \text{ Retain is Optimal }&\text{If }A<0\text{ and }B<0\end{cases}.$ (2)
In the decision rule above, for retaining to be the optimal decision, we require the expected MSE for recalibrating and refitting to be greater than retraining the model. This is highly unlikely as it would require the model refitting cost and the data collection cost to be high enough that a recalibration is worse than doing nothing. In such a scenario, the practitioner has practically no choices, so while retaining is included for completeness, we shall not see it chosen in the simulated or real data simulations below. In the Supplementary Material, we show the explicit mapping between this decision rule and the setup in the case of portolio optimization.
2.2.1 Choosing $\lambda$ and $\theta$
The above decision rule (2) relies on user-specified preferences for risk ($\lambda$) and ambiguity ($\theta$). Soliciting such preferences is an active area of research, and there exists a large body of literature on the topic (Guidolin and Rinaldi, 2012; Buhren et al., 2021). Two promising approaches, as described in Bonhomme and Weidner (2022), are a context-specific approach versus a hypothesis testing approach. The context-specific approach elicits informative choices of $\lambda$ and $\theta$ using prior knowledge of the type of misspecification. The hypothesis testing approach is most salient when a user lacks such domain knowledge.
We develop a strategy based on the hypothesis testing approach. Immediately after AI/ML model training, we expect minimal drift, and thus retaining is likely the best option. As we are dealing with a stochastic decision-making process, there is always a nonzero probability that we could choose to refit or recalibrate during this period. We can decrease the probability that refitting or recalibrating is chosen by increasing $\lambda$ and $\theta$, our aversion parameters. We choose values of $\lambda$ and $\theta$ such that, early on, we have a less than $\alpha$ probability of choosing anything but to retain the current model by setting: